# An Exploration of Default Images in Text-to-Image Generation

### v5

@@ Abstract.
copyright: rightsretainedjournalyear: 2025ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Neural networksccs: Applied computing Media arts
@@ 1. Introduction
Generative artificial intelligence (AI) has swept into mainstream. One popular type of generative AI is text-to-image (TTI) generation, where images are created by prompting the generative model with text (prompting-ai-art; palgrave; feng-etal-2023-uncovering). The prompt consists of a textual description of the envisioned image, including keywords that are known to trigger specific subjects, forms, and styles (taxonomy; ghibli; 3491102.3501825.pdf). The practice of ‘prompt engineering’ is key to TTI generation (palgrave). Practitioners craft a prompt, observe the outcome, and adjust the prompt until a satisfactory outcome is achieved (prompting-ai-art).
@@
While there is a growing number of studies investigating the practice of prompt engineering or prompt design for TTI generation (e.g., (prompting-ai-art; 3613904.3642861.pdf; 3491102.3501825.pdf; 3527927.3532792.pdf; 3706598.3713166.pdf; ICCC-2023_paper_104.pdf; 3613905.3650947.pdf)), and the design of user interfaces for prompt engineering (e.g., (Promptify; 3613904.3642016.pdf; 2508.06065.pdf; GenAssist)), less attention has been given to how models behave when prompts are misunderstood or ambiguously interpreted. TTI models are trained to always yield an output, even if the input prompt contains terms and concepts not known to the model. When such a prompt is given to the TTI model, it may generate what we call default images – visually similar images resulting from dissimilar prompts (see figures 1 and 5, and Appendix 10). These images are seemingly unrelated to the input prompt and, therefore, often do not match the user’s expectations.
@@
Default images reveal gaps in the generative model’s vocabulary and capabilities. For instance, a creative practitioner may prompt a TTI model with the Tagalog word “pagpag” (referring to a dish made from leftover food scavenged from garbage dumps), resulting in a default image (e.g., a painting of a woman with a cloud covering her face). This indicates that the prompt was not understood and the practitioner should revise the wording to improve the generated image. Default images can put an early end to a potentially unfruitful pathway in the practitioner’s prompting of the TTI model, but they can also derail the entire creative process, leading the user to abandon the quest to generate their desired image (cf. Section 2.2). Therefore, default images and their recognition play a crucial role in the creative practice of prompt engineering and for discovering the limits of TTI models and their (computational) creativity. Understanding how generative systems respond to unfamiliar or ill-formed prompts is essential for designing interactive systems that support creative exploration and for providing feedback to users. In human–AI co-creation, unexpected outputs (such as default images) can disrupt the feedback loop between user intent and system output, introducing ambiguity or undifferentiation. This undermines trust, perceived accuracy, usability, and user agency (Kreminski2025; Honeycutt202063; Riveiro2021). The misalignment between user intent and outputs can lead to frustration and a sense of lost control over the creative process (Louie2020). Investigating the phenomenon of default images contributes to broader HCI discussions on transparency, interpretability, and the design of meaningful interactions with generative models. Default images have not received attention in prior research.
@@
  • Open-source contribution: We share a dataset with 189,431 prompts and 757,724 candidates of default images, enabling future research on default images. [^2] Link after acceptance.
@@ 2.1. What are Default Images?
Some inputs to TTI generation systems may seemingly fail to generate the intended results from the creative practitioner’s perspective. We call this failure mode default images. We use the term ‘default’ to denote that TTI models resort to generating these images when they do not recognize their input. This could be because the input terms were not part of the generative model’s training dataset or because the terms in the prompt do not invoke strong imagery. In these cases, distinct prompts may yield near-identical images. In machine learning, a related concept is called mode collapse (NIPS-2014-generative-adversarial-nets-Paper.pdf), a failure that limits the usefulness of generative models for image synthesis.
The defining characteristic of default images is that they are encountered in response to multiple different and unrelated prompts. We demonstrate this in Figure 1 and Figure 5. While there are often slight variations to the default images, they can be remarkably stable in terms of motif (i.e., subject) and style (cf. Figure 5), making default images distinctly recognizable. Experienced practitioners and in particular creative practitioners who seek to deviate from common prompts will likely have encountered default images in their prompting. This paradoxical situation, which may go unnoticed by most users, reflects the current limitations of TTI models. As MANOVICH observes, AI image generators “deal well with frequently occurring items and patterns in the data but do not know what to do with the infrequent and uncommon” (MANOVICH). These models “are much better at generating visuals of popular subjects and aesthetics than unfamiliar or rare ones.”
Artists and other creative practitioners often seek to generate images that are novel (ko2023large), a key component of creativity (Runco01012012). Default images may pose a particular challenge to these users, since default images with their generic visuals tend to occur with uncommon and highly original prompts. Default images may satisfy users with less specific image requirements or lower task persistence. In these cases, users may engage in cognitive gap-filling processes (wertheimer1923gestalt) or construct personal associations (bartlett1932remembering), resulting in images that appear more relevant to the prompts than they actually are. This effect may be further amplified when the generated images have strong aesthetic appeal, since aesthetically pleasing stimuli are often judged more favorably (reber2004processing) and attractive designs are often perceived as more effective, regardless of their actual function (kurosu1995apparent). Users may accept visually appealing default outputs rather than investing effort to generate better alternatives. It is detrimental to creativity overall if TTI users are pushed towards repetitive, generic images in response to their novel prompting. This issue makes the phenomenon of default images particularly relevant to visual creativity and originality.
@@ 2.2. Importance of Studying Default Images
Many Midjourney users type long, descriptive prompts containing specific keywords, representing known concepts and style modifiers (taxonomy; prompting-ai-art). Known concepts reliably elicit the desired imagery, but they also act like ‘magnets’ that overshadow any unknown concepts, moving the output away from default images towards imagery related to the known concept. For instance, including “trump” or even “donald” in a Midjourney prompt may invoke images of a certain U.S. president in many different contexts. This means the majority of Midjourney users will likely not encounter default images in their image creation efforts. So why study default images?
@@
Understanding default images informs data and model quality improvements. Default images can indicate that the training data lacks diversity or contains biased patterns, for instance, highlighting insufficient coverage of low-resource languages, such as Tagalog, an Austronesian language spoken as a first language by about a quarter of the population of the Philippines. Addressing these cases can enhance dataset curation and fine-tuning, promoting more inclusive and culturally diverse images. Studying this phenomenon can help monitor the risk of default images entering training datasets, which could amplify the self-reinforcing loop of AI-generated media (ARIELLI; 2505.08803.pdf; 2407.17493.pdf).
@@ 3.1. Text-to-Image Generation and Prompt Engineering
While TTI systems attempt to match human expectations, the prompting language is loose and it is near impossible to describe every detail of an envisioned output in one textual prompt, with a limited number of characters (Gafni:2022). Further, the model may lack relevant prior knowledge and context. While special prompt keywords can address some of these shortcomings, such as improving the resulting image quality and imbuing it with certain styles (taxonomy), TTI models have specific failure modes that we review in the following section.
@@ 3.2. Failure Modes in Generative Models
TTI outputs frequently diverge from user intent due to the limitations of prompting as an interface (Morris). One reason is semantic ambiguity. Polysemous terms (such as “bank”, “bat”, or “ball”) can yield images that reflect unintended word senses, creating a mismatch between the user’s mental model and the generative model’s output (white2022schrodingersbatdiffusionmodels; 10.1016/j.ijhcs.2024.103375). Users may intentionally use polysemous and ambigous terms in their creative practice. So-called “magic terms” – lexical additives intentionally designed to induce randomness or probe latent failure boundaries – may result in unpredictable or uncanny imagery (taxonomy; maldaner2025miragemultimodelinterfacereviewing; 2403.12075.pdf).
@@
3491102.3501825.pdf presented a manual analysis of failure modes in text-to-image models (3491102.3501825.pdf), identifying recurring breakdowns including underspecified prompts, overloaded terms, and model over-reliance on stereotypical associations. Their findings highlight that many failures stem from subtle mismatches between user phrasing and model priors. This research was later extended to image prompts. 3527927.3532792.pdf categorized sources of model misinterpretation and proposed design strategies to expose latent failure cases (3527927.3532792.pdf). These studies contribute to a broader understanding of how prompt formulation interacts with failure patterns in TTI systems. Failure modes are relevant in interfaces for reviewing and auditing TTI models (maldaner2025miragemultimodelinterfacereviewing; 2507.17922.pdf). 2403.12075.pdf presented Adversarial Nibbler, a system for crowdsourcing the collection of adversarial examples (2403.12075.pdf), and HCOMP_24_WIP_4.pdf presented Mirage, a system for crowdsourced auditing of generative models (Text-35672-1-2-20241013.pdf; HCOMP_24_WIP_4.pdf).
@@
Beyond adversarial prompts, training artifacts can manifest as bizarre text–image associations. As these pairs have been drawn from the internet, they can include incorrect assignments or mismatched correspondences (feng-etal-2023-uncovering). daras2022 found “strange” associations between special terms in prompts and concepts in generated images (daras2022). They speculated that these associations in TTI models constituted a “secret language”. williams2024drawlunderstandingeffectsnonmainstream investigated the effects of non-mainstream dialects in TTI (williams2024drawlunderstandingeffectsnonmainstream), finding that minimal changes in prompts (including dialect) can manifest biases in TTI models. Community reports of special tokens in language models (and by extension, in TTI models), such as “SolidGoldMagikarp”, highlight spontaneous ‘glitch’ phenomena that can arise without adversarial intent (LessWrong; SolidGoldMagikarp).
@@
In addition to semantic and training-related failures, TTI models exhibit high stochasticity, introducing a form of variability that itself constitutes a failure mode. Identical prompts can produce substantially different images across runs, complicating both user expectation and scientific reproducibility. Strategies such as sampling multiple random seeds and aggregating outputs have been proposed to mitigate this variability (Liu2022), and studies show that seed choice alone can influence the emergence of rare or novel concepts (samuel2023generatingimagesrareconcepts). Failures of prompt recognition – where the model fails to parse novel or unsupported inputs and defaults to placeholder imagery – remain underexamined. Our work targets this gap by presenting a systematic approach for eliciting, identifying, and classifying default-image failures in Midjourney.
@@ 3.3. Detecting Mode Collapse in Text-to-Image Models
Mode collapse is a failure mode in generative models where the model produces less diverse outputs than expected. It was first noted by NIPS-2014-generative-adversarial-nets-Paper.pdf in Generative Adversarial Networks (GANs) (NIPS-2014-generative-adversarial-nets-Paper.pdf). Default images share characteristics with mode collapse: both involve gravitating toward a limited set of outputs and occur only in response to specific prompts.
Various techniques for detecting and mitigating mode collapse in TTI models have been developed and applied to GANs and diffusion models. 978-3-031-90341-0_20 studied early detection of mode collapse through loss monitoring (978-3-031-90341-0_20). Other techniques include multi-generator approaches (e.g., MTC-GAN (Zhang20217789)), regularization techniques (e.g., ICCR (10315510)), contrastive learning (e.g., MCL (s40747-022-00924-1.pdf)), and novel architectures (e.g., style-guided discriminator (Choi_Kim_Song_2022)). 1910.11626.pdf used semantic segmentation networks to compare the distribution of objects in generated images with the target distribution to identify omitted classes (1910.11626.pdf). 10334839 developed Connective Novelty Detection, a method that uses an autoencoder and a binary classifier to detect out-of-distribution samples. By creating negative instances close to the positive class distribution and combining real and generated samples, the method improves the robustness of detecting mode collapse (10334839). Gong2023 used Feature Distribution Divergence Minimization to minimize the divergence between the distributions of real and generated features, reducing the learning pressure on the discriminator and addressing mode collapse without requiring prior knowledge or manual design (Gong2023). Duym_2025_WACV developed the Mode Collapse Entropy (MCE) score (Duym_2025_WACV) to overcome weaknesses of previous metrics in real-world scenarios.
@@
Several inference-time methods have been proposed for detecting mode collapse in black-box models. Statistical, sampling-based tools can be used to visualize and quantify intra-mode collapse without access to training data or model parameters, providing a practical approach for black-box diagnosis and calibration (Wu2021). Inversion-based inference perturbation (IIP) alters the inference process to detect memorized images, using unconditional DDIM inversion to extract latent codes and identify images exhibiting collapse-like behavior (Jiang202590485). EncoderMI leverages membership inference against contrastive learning–based image encoders, exploiting their overfitting tendencies to reveal whether inputs belong to the training set, which can indirectly indicate mode collapse (Liu20212081). Binary noise guidance learning (BnGLGAN) introduces distribution-guided noise reconstruction to stabilize image translation and mitigate collapse (Zhang2024).

### v6

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790681isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Neural networksccs: Applied computing Media arts
@@ 1. Introduction
Generative artificial intelligence (AI) has swept into mainstream. One popular type of generative AI is text-to-image (TTI) generation, where images are created by prompting the generative model with text (Oppenlaender et al., 2024; Oppenlaender, 2024; Feng et al., 2023). The prompt consists of a textual description of the envisioned image, including keywords that are known to trigger specific subjects, forms, and styles (Oppenlaender, 2023; Placido, 2025; Liu and Chilton, 2022b). The practice of ‘prompt engineering’ is key to TTI generation (Oppenlaender, 2024). Practitioners craft a prompt, observe the outcome, and adjust the prompt until a satisfactory outcome is achieved (Oppenlaender et al., 2024).
@@
While there is a growing number of studies investigating the practice of prompt engineering or prompt design for TTI generation (e.g., (Oppenlaender et al., 2024; Mahdavi Goloujeh et al., 2024a; Liu and Chilton, 2022b; Qiao et al., 2022; Subramonyam et al., 2025; Bird, 2023; Mahdavi Goloujeh et al., 2024c)), and the design of user interfaces for prompt engineering (e.g., (Mahdavi Goloujeh et al., 2024b; Arawjo et al., 2024; Lee et al., 2025; Huh et al., 2023)), less attention has been given to how models behave when prompts are misunderstood or ambiguously interpreted. TTI models are trained to always yield an output, even if the input prompt contains terms and concepts not known to the model. When such a prompt is given to the TTI model, it may generate what we call default images – visually similar images resulting from dissimilar prompts (see figures 1 and 5, and Appendix 10). These images are seemingly unrelated to the input prompt and, therefore, often do not match the user’s expectations.
@@
Default images reveal gaps in the generative model’s vocabulary and capabilities. For instance, a creative practitioner may prompt a TTI model with the Tagalog word “pagpag” (referring to a dish made from leftover food scavenged from garbage dumps), resulting in a default image (e.g., a painting of a woman with a cloud covering her face). This indicates that the prompt was not understood and the practitioner should revise the wording to improve the generated image. Default images can put an early end to a potentially unfruitful pathway in the practitioner’s prompting of the TTI model, but they can also derail the entire creative process, leading the user to abandon the quest to generate their desired image (cf. Section 2.2). Therefore, default images and their recognition play a crucial role in the creative practice of prompt engineering and for discovering the limits of TTI models and their (computational) creativity. Understanding how generative systems respond to unfamiliar or ill-formed prompts is essential for designing interactive systems that support creative exploration and for providing feedback to users. In human–AI co-creation, unexpected outputs (such as default images) can disrupt the feedback loop between user intent and system output, introducing ambiguity or undifferentiation. This undermines trust, perceived accuracy, usability, and user agency (Kreminski, 2025; Honeycutt et al., 2020; Riveiro and Thill, 2021). The misalignment between user intent and outputs can lead to frustration and a sense of lost control over the creative process (Louie et al., 2020). Investigating the phenomenon of default images contributes to broader HCI discussions on transparency, interpretability, and the design of meaningful interactions with generative models. Default images have not received attention in prior research.
@@
  • Open-source contribution: We share a dataset with 189,432 prompts and 757,728 candidates of default images, enabling future research on default images (see https://huggingface.co/ datasets/tti-dev/default-images).
@@ 2.1. What are Default Images?
Some inputs to TTI generation systems may seemingly fail to generate the intended results from the creative practitioner’s perspective. We call this failure mode default images. We use the term ‘default’ to denote that TTI models resort to generating these images when they do not recognize their input. This could be because the input terms were not part of the generative model’s training dataset or because the terms in the prompt do not invoke strong imagery. In these cases, distinct prompts may yield near-identical images. In machine learning, a related concept is called mode collapse (Goodfellow et al., 2014), a failure that limits the usefulness of generative models for image synthesis.
The defining characteristic of default images is that they are encountered in response to multiple different and unrelated prompts. We demonstrate this in Figure 1 and Figure 5. While there are often slight variations to the default images, they can be remarkably stable in terms of motif (i.e., subject) and style (cf. Figure 5), making default images distinctly recognizable. Experienced practitioners and in particular creative practitioners who seek to deviate from common prompts will likely have encountered default images in their prompting. This paradoxical situation, which may go unnoticed by most users, reflects the current limitations of TTI models. As Manovich observes, AI image generators “deal well with frequently occurring items and patterns in the data but do not know what to do with the infrequent and uncommon” (Manovich, 2024). These models “are much better at generating visuals of popular subjects and aesthetics than unfamiliar or rare ones.”
Artists and other creative practitioners often seek to generate images that are novel (Ko et al., 2023), a key component of creativity (Runco and Jaeger, 2012). Default images may pose a particular challenge to these users, since default images with their generic visuals tend to occur with uncommon and highly original prompts. Default images may satisfy users with less specific image requirements or lower task persistence. In these cases, users may engage in cognitive gap-filling processes (Wertheimer, 1938) or construct personal associations (Bartlett, 1995), resulting in images that appear more relevant to the prompts than they actually are. This effect may be further amplified when the generated images have strong aesthetic appeal, since aesthetically pleasing stimuli are often judged more favorably (Reber et al., 2004) and attractive designs are often perceived as more effective, regardless of their actual function (Kurosu and Kashimura, 1995). Users may accept visually appealing default outputs rather than investing effort to generate better alternatives. It is detrimental to creativity overall if TTI users are pushed towards repetitive, generic images in response to their novel prompting. This issue makes the phenomenon of default images particularly relevant to visual creativity and originality.
@@ 2.2. Importance of Studying Default Images
Many Midjourney users type long, descriptive prompts containing specific keywords, representing known concepts and style modifiers (Oppenlaender, 2023; Oppenlaender et al., 2024). Known concepts reliably elicit the desired imagery, but they also act like ‘magnets’ that overshadow any unknown concepts, moving the output away from default images towards imagery related to the known concept. For instance, including “trump” or even “donald” in a Midjourney prompt may invoke images of a certain U.S. president in many different contexts. This means the majority of Midjourney users will likely not encounter default images in their image creation efforts. So why study default images?
@@
Understanding default images informs data and model quality improvements. Default images can indicate that the training data lacks diversity or contains biased patterns, for instance, highlighting insufficient coverage of low-resource languages, such as Tagalog, an Austronesian language spoken as a first language by about a quarter of the population of the Philippines. Addressing these cases can enhance dataset curation and fine-tuning, promoting more inclusive and culturally diverse images. Studying this phenomenon can help monitor the risk of default images entering training datasets, which could amplify the self-reinforcing loop of AI-generated media (Arielli, 2024; Hu et al., 2025; Yoon et al., 2025).
@@ 3.1. Text-to-Image Generation and Prompt Engineering
While TTI systems attempt to match human expectations, the prompting language is loose and it is near impossible to describe every detail of an envisioned output in one textual prompt, with a limited number of characters (Gafni et al., 2022). Further, the model may lack relevant prior knowledge and context. While special prompt keywords can address some of these shortcomings, such as improving the resulting image quality and imbuing it with certain styles (Oppenlaender, 2023), TTI models have specific failure modes that we review in the following section.
@@ 3.2. Failure Modes in Generative Models
TTI outputs frequently diverge from user intent due to the limitations of prompting as an interface (Morris, 2024). One reason is semantic ambiguity. Polysemous terms (such as “bank”, “bat”, or “ball”) can yield images that reflect unintended word senses, creating a mismatch between the user’s mental model and the generative model’s output (White and Cotterell, 2022; Rapp et al., 2025). Users may intentionally use polysemous and ambigous terms in their creative practice. So-called “magic terms” – lexical additives intentionally designed to induce randomness or probe latent failure boundaries – may result in unpredictable or uncanny imagery (Oppenlaender, 2023; Maldaner et al., 2025; Quaye et al., 2024).
@@
Liu and Chilton presented a manual analysis of failure modes in text-to-image models (Liu and Chilton, 2022b), identifying recurring breakdowns including underspecified prompts, overloaded terms, and model over-reliance on stereotypical associations. Their findings highlight that many failures stem from subtle mismatches between user phrasing and model priors. This research was later extended to image prompts. Qiao et al. categorized sources of model misinterpretation and proposed design strategies to expose latent failure cases (Qiao et al., 2022). These studies contribute to a broader understanding of how prompt formulation interacts with failure patterns in TTI systems. Failure modes are relevant in interfaces for reviewing and auditing TTI models (Maldaner et al., 2025; Quaye et al., 2025). Quaye et al. presented Adversarial Nibbler, a system for crowdsourcing the collection of adversarial examples (Quaye et al., 2024), and Maldaner et al. presented Mirage, a system for crowdsourced auditing of generative models (Deng et al., 2024; Maldaner et al., 2025).
@@
Beyond adversarial prompts, training artifacts can manifest as bizarre text–image associations. As these pairs have been drawn from the internet, they can include incorrect assignments or mismatched correspondences (Feng et al., 2023). Daras and Dimakis found “strange” associations between special terms in prompts and concepts in generated images (Daras and Dimakis, 2022). They speculated that these associations in TTI models constituted a “secret language”. Williams et al. investigated the effects of non-mainstream dialects in TTI (Williams et al., 2024), finding that minimal changes in prompts (including dialect) can manifest biases in TTI models. Community reports of special tokens in language models (and by extension, in TTI models), such as “SolidGoldMagikarp”, highlight spontaneous ‘glitch’ phenomena that can arise without adversarial intent (LessWrong, n.d.; Rumbelow and Watkins, 2023).
@@
In addition to semantic and training-related failures, TTI models exhibit high stochasticity, introducing a form of variability that itself constitutes a failure mode. Identical prompts can produce substantially different images across runs, complicating both user expectation and scientific reproducibility. Strategies such as sampling multiple random seeds and aggregating outputs have been proposed to mitigate this variability (Liu and Chilton, 2022a), and studies show that seed choice alone can influence the emergence of rare or novel concepts (Samuel et al., 2024). Failures of prompt recognition – where the model fails to parse novel or unsupported inputs and defaults to placeholder imagery – remain underexamined. Our work targets this gap by presenting a systematic approach for eliciting, identifying, and classifying default-image failures in Midjourney.
@@ 3.3. Detecting Mode Collapse in Text-to-Image Models
Mode collapse is a failure mode in generative models where the model produces less diverse outputs than expected. It was first noted by Goodfellow et al. in Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Default images share characteristics with mode collapse: both involve gravitating toward a limited set of outputs and occur only in response to specific prompts.
Various techniques for detecting and mitigating mode collapse in TTI models have been developed and applied to GANs and diffusion models. Barsha and Eberle studied early detection of mode collapse through loss monitoring (Barsha and Eberle, 2025). Other techniques include multi-generator approaches (e.g., MTC-GAN (Zhang et al., 2021)), regularization techniques (e.g., ICCR (Tominaga and Seo, 2023)), contrastive learning (e.g., MCL (Gou et al., 2023)), and novel architectures (e.g., style-guided discriminator (Choi et al., 2022)). Bau et al. used semantic segmentation networks to compare the distribution of objects in generated images with the target distribution to identify omitted classes (Bau et al., 2019). Hashemi et al. developed Connective Novelty Detection, a method that uses an autoencoder and a binary classifier to detect out-of-distribution samples. By creating negative instances close to the positive class distribution and combining real and generated samples, the method improves the robustness of detecting mode collapse (Hashemi et al., 2023). Gong et al. used Feature Distribution Divergence Minimization to minimize the divergence between the distributions of real and generated features, reducing the learning pressure on the discriminator and addressing mode collapse without requiring prior knowledge or manual design (Gong et al., 2023). Duym et al. developed the Mode Collapse Entropy (MCE) score (Duym et al., 2025) to overcome weaknesses of previous metrics in real-world scenarios.
@@
Several inference-time methods have been proposed for detecting mode collapse in black-box models. Statistical, sampling-based tools can be used to visualize and quantify intra-mode collapse without access to training data or model parameters, providing a practical approach for black-box diagnosis and calibration (Wu et al., 2021). Inversion-based inference perturbation (IIP) alters the inference process to detect memorized images, using unconditional DDIM inversion to extract latent codes and identify images exhibiting collapse-like behavior (Jiang et al., 2025). EncoderMI leverages membership inference against contrastive learning–based image encoders, exploiting their overfitting tendencies to reveal whether inputs belong to the training set, which can indirectly indicate mode collapse (Liu et al., 2021). Binary noise guidance learning (BnGLGAN) introduces distribution-guided noise reconstruction to stabilize image translation and mitigate collapse (Zhang et al., 2024).